OpenStudy (anonymous):

how can the division property of equality be used to solve multiplication equations

OpenStudy (unklerhaukus):

5x = 10 if we divide both sides of the equation by 5 5x/5= 10/5 then on the left hand side the 5's will cancel x = 10/5 simplifying x = 2